# Optimistic Estimation of Convergence in Markov Chains with the Average-Mixing Time

### v2

@@ 4   Implications at various scales of state space
In infinitely denumerable state spaces, recall that for any $s\in\mathbb{N}$, the well-definition of the estimation rate for the $\beta(s)$-mixing coefficient (Theorem 3.1) hinges on the finiteness of the following quantity
$\mathcal{J}^{(s)}_{p}\doteq\left\|Q^{(s)}\right\|_{1-1/p}^{1-1/p},$
where $p\in[1,\infty)$ is a user-fixed parameter. It is instructive to observe that on one hand
$\lim_{p\to\infty}\mathcal{J}_{p}^{(s)}=1,$
while on the other hand, $\mathcal{B}_{p}^{(s)}$ increases with $p$,
$\lim_{p\to\infty}\mathcal{B}_{p}^{(s)}\asymp\frac{n}{s},$
leading to a trade-off in the choice of the parameter in Theorem 3.1. In particular, provided a control which holds uniformly over $s$,
$\mathcal{J}_{p,\xi}\doteq\sup\left\{\mathcal{J}^{(s)}_{p}\colon s\in\mathbb{N},s\leq t^{\sharp}_{\mathsf{mix}}(\xi)\right\},$ (7)
the free parameter $p$ can be appropriately selected in Theorem 3.4, in order to optimize the upper bound on the sample complexity of estimating the average-mixing time. It is also noteworthy that since the $\ell_{1-1/p}$ quasi-norm and the $(1-1/p)$th order Rényi entropy $H_{1-1/p}$ of a distribution $\mu\in\mathcal{P}(\mathcal{X})$ are related by the expression
$p\log\left\|\mu\right\|_{1-1/p}^{1-1/p}=H_{1-1/p}(\mu),$
the quantity $\mathcal{J}_{p,\xi}$ pertains to a specific entropic property inherent in the Markov process. Unfortunately, this term is generally not easily amenable to analysis; hence, our focus lies in establishing bounds for its evaluation. Under ergodicity of $P$, the pair-wise limit
$\lim_{s\to\infty}Q^{(s)}(x,x^{\prime})=\pi(x)\pi(x^{\prime}),\qquad\forall x,x^{\prime}\in\mathcal{X},$
hints at a heuristic lower bound in $\left\|\pi\right\|_{1-1/p}^{2(1-1/p)}$ on $\mathcal{J}_{p,\xi}$. Using the stationary regime as a baseline, it seems natural to examine chains where the entropic term does not undergo relative explosive growth in the transient regime. For instance, for some multiplicative factor $K\geq 1$, an upper bound
$\mathcal{J}_{p}^{(s)}\leq K\left\|\pi\right\|^{2(1-1/p)}_{1-1/p},$ (8)
with a controlled $K$ readily simplifies the sample complexities appearing in Section 3. In the present section, we proceed to discuss bounds on the entropic quantities $\mathcal{J}_{p}^{(s)}$ and $\mathcal{J}_{p,\xi}$ in terms of state-wise conditional entropy, structural properties of the connection graph of the Markov chain, or the alphabet size. We conclude with a discussion on the binary state space, where the dominant complexity stems from the mixing bottlenecks, revealing that our results hold implications at every scale.
@@ 4.1   Ergodic properties
In this subsection, we derive various controls of $\mathcal{J}_{p}^{(s)}$ based on ergodic properties of the chain. Specifically, we introduce three conditions. Condition 1 is the classical notion of $V$-geometric ergodicity, which is known to be weaker than uniform ergodicity. Condition 2 imposes a stronger criterion, under which an ergodic chain is necessarily uniformly ergodic. Both Conditions 1 and 2 allow to derive bounds on $\mathcal{J}_{p}^{(s)}$ that are uniform over $s$. However, these conditions are quite stringent. We introduce a natural extension of these two conditions, Condition 3, which also allows to control $\mathcal{J}_{p}^{(s)}$. However, under Condition 3, this bound might be growing (sub-linearly) with $s$. We summarize the relations between these conditions in Figure 1.
Bound on ${\text{Bound on }\mathcal{J}_{p}^{(s)}\text{ uniform over }s}$ Condition 1Condition 2Condition 3Bound on ${\text{Bound on }\mathcal{J}_{p}^{(s)}\text{ sub-linear in }s}$Corollary 4.1Corollary 4.1Lemma 4.2Lemma 4.3Lemma 4.1
Figure 1: Relation between the various ergodic conditions in Subsection 4.1. In this diagram, $\Rightarrow$ is to be read as ”implies under the assumptions of the result referred to”.
Condition 1 ($V$-geometric ergodicity (Douc et al., 2018, Section 18)).
Let $P\in\mathcal{W}(\mathcal{X})$ be a $\pi$-stationary transition operator. We say that $P$ is $V$-geometrically ergodic for some function $V\colon\mathcal{X}\to[1,\infty)$, when there exist constants $\theta\in(0,1)$ and $M\in\mathbb{R}_{+}$ such that for any $x\in\mathcal{X}$,
$\left\|e_{x}P^{s}-\pi\right\|_{V}\leq MV(x)\theta^{s},$
where for any $\nu\in\mathbb{R}^{\mathcal{X}}$, we wrote
$\left\|\nu\right\|_{V}\doteq\sup_{x\in\mathcal{X}}V(x)\nu(x).$
Condition 2 (Existence of a dominating distribution).
Let $P\in\mathcal{W}(\mathcal{X})$ be a $\pi$-stationary transition operator. Suppose that there exists a universal constant $C\in\mathbb{R}_{+}$ and a probability measure $\nu\in\mathcal{P}(\mathcal{X})$, such that for all $x^{\prime}\in\mathcal{X}$,
$\sup_{x\in\mathcal{X}}P(x,x^{\prime})\leq C\nu(x^{\prime}).$
Condition 3 (Weak $V$-geometric ergodicity).
@@
Let $P\in\mathcal{W}(\mathcal{X})$ be a $\pi$-stationary transition operator. We say that $P$ is weakly $V$-geometrically ergodic when there exists functions
$V\colon\mathcal{X}\to[1,\infty),\qquad D\colon\mathcal{X}\to[1,\infty),$
and a constant $b\geq 0$ such that for any $s\in\mathbb{N}$,
$\sup_{x^{\prime}\in\mathcal{X}}V(x^{\prime})(P^{s}(x,x^{\prime})-\pi(x^{\prime}))\leq D(x)+b(s-1).$
Lemma 4.1.
Let $P\in\mathcal{W}(\mathcal{X})$ be a $\pi$-stationary weakly $V$-geometrically ergodic transition operator, with $V,D,b$ defined as in Condition 3, and let $p\geq 1$ such that
$\left\|1/V\right\|_{1-1/p}^{1-1/p}<\infty,\qquad\left\|\pi\circ D\right\|_{1-1/p}^{1-1/p}<\infty.$
Then,
$\mathcal{J}_{p}^{(s)}\leq\left\|\pi\right\|_{1-1/p}^{2(1-1/p)}+\left\|1/V\right\|_{1-1/p}^{1-1/p}\left(\left\|\pi\circ D\right\|_{1-1/p}^{1-1/p}+\left\|\pi\right\|_{1-1/p}^{1-1/p}[b(s-1)]^{1-1/p}\right).$
Proof.
Directly from the assumptions,
$\mathcal{J}_{p}^{(s)}$ $\leq\sum_{x,x^{\prime}\in\mathcal{X}}\left(\pi(x)\left[\pi(x^{\prime})+\frac{1}{V(x^{\prime})}\left(D(x)+b(s-1)\right)\right]\right)^{1-1/p}$
$\leq\left\|\pi\right\|_{1-1/p}^{2(1-1/p)}+\sum_{x,x^{\prime}\in\mathcal{X}}\left[\pi(x)\frac{1}{V(x^{\prime})}\left(D(x)+b(s-1)\right)\right]^{1-1/p}$
$\leq\left\|\pi\right\|_{1-1/p}^{2(1-1/p)}+\left\|1/V\right\|_{1-1/p}^{1-1/p}\left(\left\|\pi\circ D\right\|_{1-1/p}^{1-1/p}+\left\|\pi\right\|_{1-1/p}^{1-1/p}[b(s-1)]^{1-1/p}\right),$
where the two inequalities stem from the sub-additivity of the function $x\mapsto x^{1-1/p}$. ∎
Corollary 4.1 (Under $V$-geometric ergodicity).
Let $P$ a $\pi$-stationary $V$-geometrically ergodic transition operator, with $V,M,\theta$ defined as in Condition 1. Then Condition 3 is satisfied with
$D=M\theta V\text{ and }b=0.$
If the following summability conditions are met for some $p\geq 1$,
$\left\|1/V\right\|_{1-1/p}^{1-1/p}<\infty,\qquad\left\|\pi\circ V\right\|_{1-1/p}^{1-1/p}<\infty,$
we obtain a control of $\mathcal{J}_{p}^{(s)}$ uniformly over $s$:
$\sup_{s\in\mathbb{N}}\mathcal{J}_{p}^{(s)}\\ \leq\|\pi\|_{1-1/p}^{2(1-1/p)}+\left\|1/V\right\|_{1-1/p}^{1-1/p}\left\|\pi\circ D\right\|^{1-1/p}_{1-1/p}.$
Proof.
We first check Condition 3. For any $x,x^{\prime}\in\mathcal{X}$,
$V(x)(P^{s}(x,x^{\prime})-\pi(x^{\prime}))$ $\leq\sup_{x^{\prime\prime}\in\mathcal{X}}V(x^{\prime\prime})|P^{s}(x,x^{\prime\prime})-\pi(x^{\prime\prime})|$
$=\left\|e_{x}P^{s}-\pi\right\|_{V}\leq MV(x)\theta^{s}\leq MV(x)\theta,$
where the second inequality stems from the definition of $V$-geometric ergodicity. Thus, Condition 3 is satisfied with $D=M\theta V$ and $b=0$. An application of Lemma 4.1 gives the the second claim. ∎
@@
Under the assumption $D\propto V$, we can clarify the summability assumptions in Corollary 4.1.
  • Let us first assume that $\pi(x)\propto x^{-\alpha}$ for $\alpha\in(1,\infty)$. Then, for $V(x)\propto D(x)\propto x^{\beta}$, with $\beta\in\mathbb{R}_{+}$, $\left\|1/V\right\|_{1-1/p}^{1-1/p}<\infty$ holds only when $\beta>p/(p-1)$, and $\left\|\pi\circ D\right\|_{1-1/p}^{1-1/p}<\infty$ holds only when $\alpha-\beta>p/(p-1)$, thus we must have $p/(p-1)<\beta<\alpha-p/(p-1)$, which will only be possible for $\alpha>2p/(p-1)$.
  • Let us now assume now that $\pi(x)\propto a^{-x}$ for $a\in(0,1)$. Then, for $V(x)\propto D(x)\propto x^{\beta}$, $\left\|\pi\circ D\right\|_{1-1/p}^{1-1/p}<\infty$ always holds but we still require $\beta>p/(p-1)$ in order for $\left\|1/V\right\|_{1-1/p}^{1-1/p}<\infty$ to hold.
Lemma 4.2 (Under the existence of a dominating distribution).
Assume Condition 2 is satisfied. If moreover, for $p\geq 1$,
$\left\|\pi\right\|_{1-1/p}^{1-1/p}<\infty,\qquad\left\|\nu\right\|_{1-1/p}^{1-1/p}<\infty,$
then,
$\sup_{s\in\mathbb{N}}\mathcal{J}_{p}^{(s)}\leq C^{1-1/p}\left\|\nu\right\|_{1-1/p}^{1-1/p}\left\|\pi\right\|_{1-1/p}^{1-1/p}.$
Proof.
It directly follows from the assumptions that for any $x,x^{\prime}\in\mathcal{X}$, and for any $s\in\mathbb{N}$,
$\begin{split}Q^{(s)}(x,x^{\prime})&=\pi(x)\sum_{x^{\prime\prime}\in\mathcal{X}}P^{s-1}(x,x^{\prime\prime})P(x^{\prime\prime},x^{\prime})\\ &\leq C\pi(x)\nu(x^{\prime})\sum_{x^{\prime\prime}\in\mathcal{X}}P^{s-1}(x,x^{\prime\prime})=C\pi(x)\nu(x^{\prime}).\end{split}$
As a result,
$\sum_{x,x^{\prime}\in\mathcal{X}}Q^{(s)}(x,x^{\prime})^{1-1/p}\leq C^{1-1/p}\sum_{x\in\mathcal{X}}\pi(x)^{1-1/p}\sum_{x^{\prime}\in\mathcal{X}}\nu(x^{\prime})^{1-1/p}.$
In practice, Condition 3 can be readily verified through the following lemma.
Lemma 4.3.
Let $P\in\mathcal{W}(\mathcal{X})$ be a $\pi$-stationary transition operator. Assume there exists functions $V\colon\mathcal{X}\to[1,\infty)$, $D\colon\mathcal{X}\to[1,\infty)$ and a constant $b\geq 0$ such that
$\sup_{x^{\prime}\in\mathcal{X}}V(x^{\prime})(P(x,x^{\prime})-\pi(x^{\prime}))\leq D(x)$
and, for all $x\in\mathcal{X}$, $PD(x)\leq D(x)+b$. Then, Condition 3 holds.
In Section 6, we use Lemma 4.3 to check Condition 3 on a whole family of transition operators $(P_{T})_{T\in\mathcal{T}}$. Note that Lemma 4.3 also shows that Condition 2 immediately implies Condition 3 with $D=C$, $V=1$ and $b=0$, which completes Figure 1.
Proof.
By recursion, for any $t\in\mathbb{N}$,
$\forall x\in\mathcal{X},\,P^{t}D(x)\leq D(x)+bt.$ (9)
Indeed, (9) holds for $t=1$ by assumption. Assume it holds for $t-1$, then
$P^{t}D(x)$ $=P[P^{t-1}D](x)$
$=\sum_{x^{\prime}\in\mathcal{X}}P(x,x^{\prime})[P^{t-1}D](x^{\prime})$
$\leq\sum_{x^{\prime}\in\mathcal{X}}P(x,x^{\prime})[D(x^{\prime})+b(t-1)]$
$=PD(x)+b(t-1)$
$\leq D(x)+bt.$
Thus, for any $x^{\prime\prime}\in\mathcal{X}$,
$V(x^{\prime\prime})[P^{t}(x,x^{\prime\prime})-\pi(x^{\prime\prime})]$ $=\sum_{x^{\prime}\in\mathcal{X}}P^{t-1}(x,x^{\prime})[P(x^{\prime},x^{\prime\prime})-\pi(x^{\prime\prime})]V(x^{\prime\prime})$
$\leq\sum_{x^{\prime}\in\mathcal{X}}P^{t-1}(x,x^{\prime})D(x^{\prime})$
$=P^{t-1}D(x)$
$\leq D(x)+b(t-1),$
where the last inequality follows from (9). ∎
4.2   Uniformly bounded conditional entropy
When for each state $x\in\mathcal{X}$, the conditional distribution $e_{x}P$ has bounded entropy, it is possible to guarantee finiteness of $\mathcal{J}_{p,\xi}$.
Lemma 4.4.
Let $P\in\mathcal{W}(\mathcal{X})$ be a $\pi$-stationary transition operator. Suppose that for $p\geq 1$,
$\mathcal{I}_{p}\doteq\sup_{x\in\mathcal{X}}\left\{\left\|e_{x}P\right\|^{1-1/p}_{1-1/p}\right\}<\infty.$
Then, for any $s\in\mathbb{N}$, it holds that
$\mathcal{J}_{p}^{(s)}\leq\mathcal{I}_{p}^{s}\left\|\pi\right\|_{1-1/p}^{1-1/p}.$
Proof.
Suppose that the claim holds for $s\geq 1$. Then,
$\begin{split}\mathcal{J}_{p}^{(s+1)}&=\sum_{x,x^{\prime}\in\mathcal{X}}\pi(x)^{1-1/p}\left(\sum_{x^{\prime\prime}\in\mathcal{X}}P^{s}(x,x^{\prime\prime})P(x^{\prime\prime},x^{\prime})\right)^{1-1/p}\\ &\leq\sum_{x,x^{\prime}\in\mathcal{X}}\pi(x)^{1-1/p}\sum_{x^{\prime\prime}\in\mathcal{X}}P^{s}(x,x^{\prime\prime})^{1-1/p}P(x^{\prime\prime},x^{\prime})^{1-1/p}\\ &=\sum_{x,x^{\prime\prime}\in\mathcal{X}}\pi(x)^{1-1/p}P^{s}(x,x^{\prime\prime})^{1-1/p}\sum_{x^{\prime}\in\mathcal{X}}P(x^{\prime\prime},x^{\prime})^{1-1/p}\\ &\leq\sum_{x,x^{\prime\prime}\in\mathcal{X}}\pi(x)^{1-1/p}P^{s}(x,x^{\prime\prime})^{1-1/p}\sup_{x\in\mathcal{X}}\left\|e_{x}P\right\|_{1-1/p}^{1-1/p}\\ &\leq\mathcal{J}_{p}^{(s)}\mathcal{I}_{p}.\end{split}$
The case $s=1$ is easy to verify, thus the lemma holds by induction. ∎
While a direct consequence of Lemma 4.4 is that
$\mathcal{J}_{p,\xi}\leq\mathcal{I}_{p}^{t^{\sharp}_{\mathsf{mix}}(\xi)}\left\|\pi\right\|_{1-1/p}^{1-1/p},$
guaranteeing finiteness of $\mathcal{J}_{p,\xi}$ for any $\xi\in(0,1)$, the above approach is only practical when bounding $\mathcal{J}_{p}^{(s)}$ for restrictively small values of $s$.

### v4

@@ 4  Implications at various scales of state space
Recall (refer to Theorem 3.1, Theorem 3.2) that we require a bound on the entropic term
$\mathcal{J}_{p}^{(s)}=\left\|Q^{(s)}\right\|^{1-1/p}_{(1-1/p)/2},$
where for $s\in\mathbb{N}$ and for $x,x^{\prime}\in\mathcal{X}$, we wrote $Q^{(s)}(x,x^{\prime})=\pi(x)P^{s}(x,x^{\prime})$, and for $q\in\mathbb{R}_{+}$,
$\left\|Q^{(s)}\right\|^{q}_{q}\doteq\sum_{(x,x^{\prime})\in\mathcal{X}^{2}}Q^{(s)}(x,x^{\prime})^{q},$
in order to recover an estimation rate for the individual $\beta$-mixing coefficients, whilst upper bounds on the sample complexity for estimating the average-mixing time (Theorem 3.5) require a control which holds uniformly over all skipping rates up to the order of the average-mixing time,
$\mathcal{J}_{p,\xi}=\sup\left\{\mathcal{J}^{(s)}_{p}\colon s\in\mathbb{N},s\leq t^{\sharp}_{\mathsf{mix}}(\xi)\right\}.$ (11)
Bounds generally depend on the tuning parameter $p$, which can then be appropriately selected—for instance in Theorem 3.5—in order to optimize the upper bound on the sample complexity. Unfortunately, $\mathcal{J}_{p,\xi}$ is generally not easily amenable to analysis; hence, our focus lies in establishing bounds for its evaluation. It is not necessarily the case that $\mathcal{J}_{p}^{(s)}$ is monotonous in $s$. As illustrated in Figure 3, $\mathcal{J}_{p}^{(s)}$ can exhibit a transient regime. In this case, under ergodicity of $P$, the pair-wise limit
$\lim_{s\to\infty}Q^{(s)}(x,x^{\prime})=\pi(x)\pi(x^{\prime}),\qquad\forall(x,x^{\prime})\in\mathcal{X}^{2},$
hints at a heuristic lower bound in $\left\|\pi\right\|_{(1-1/p)/2}^{2(1-1/p)}$ on $\mathcal{J}_{p,\xi}$. Using the stationary regime as a baseline, it seems natural to examine whether the entropic term undergoes relative explosive growth in the transient regime. For instance, for some multiplicative factor $K\geq 1$, an upper bound
$\mathcal{J}_{p}^{(s)}\leq K\left\|\pi\right\|^{2(1-1/p)}_{(1-1/p)/2}$, (12)
with a controlled $K$ readily streamlines the sample complexities appearing in Section 3 in terms of properties of the stationary distribution. Let us illustrate the approach by assuming the existence of a uniformly dominating distribution over all conditional distributions. Formally, let us suppose that there exists a universal constant $C\in\mathbb{R}_{+}$ and a probability measure $\nu\in\mathcal{P}(\mathcal{X})$, such that for all $x^{\prime}\in\mathcal{X}$,
$\sup_{x\in\mathcal{X}}P(x,x^{\prime})\leq C\nu(x^{\prime}).$
It follows that
$\begin{split}Q^{(s)}(x,x^{\prime})&\leq\pi(x)\sum_{x^{\prime\prime}\in\mathcal{X}}C\nu(x^{\prime})P^{s-1}(x,x^{\prime\prime})=C\pi(x)\nu(x^{\prime}),\end{split}$
and thus
$\sup_{s\in\mathbb{N}}\mathcal{J}_{p}^{(s)}\leq C^{1-1/p}\left\|\nu\right\|_{(1-1/p)/2}^{1-1/p}\left\|\pi\right\|_{(1-1/p)/2}^{1-1/p}.$
In the rest of the section, we proceed to discuss bounds on the entropic quantities $\mathcal{J}_{p}^{(s)}$ and $\mathcal{J}_{p,\xi}$ in terms of ergodic properties of the chain, structural properties of the connection graph of the Markov chain, or the alphabet size. We conclude with a discussion on the binary state space, where the dominant complexity stems from the mixing bottlenecks, revealing that our approach holds implications at every scale.
[…]
Figure 3: Transient regime for the entropic term $\mathcal{J}_{p}^{(s)}$.
@@ 4.1  Ergodic properties
In this subsection, we derive a control of $\mathcal{J}_{p}^{(s)}$ based on a decomposition and natural ergodic properties of the chain that we introduce. Let us first decompose $\sqrt{\mathcal{J}_{p}^{(s)}}$ in terms of a stationary quantity and some measure of distance to stationarity. Recall that for $a,b>0,q\in(0,1)$, it holds that $(a+b)^{q}\leq a^{q}+b^{q}$, thus
$\sqrt{\mathcal{J}_{p}^{(s)}}$ $\leq\beta_{(1-1/p)/2}(s)+\left\|\pi\right\|_{(1-1/p)/2}^{1-1/p},$
where for $q\in(0,1)$, $\beta_{q}(s)\geq\beta(s)$ is a generalization of the $\beta$-mixing coefficient defined as
[…]
Moving forward, for a function $V\colon\mathcal{X}\to[1,\infty)$ we introduce the notion of the $V_{q}$-norm of a signed measure as follows,
$\left\|\mu\right\|_{q,V}^{q}\doteq\sum_{x\in\mathcal{X}}\left|\mu(x)\right|^{q}V(x),$
@@
and we say that $P$ is $V_{q}$-geometrically ergodic when there exists $\rho_{q,V}\in(0,1)$ such that for all $s\in\mathbb{N}$,
$\sup_{x\in\mathcal{X}}\frac{\left\|e_{x}P^{s}-\pi\right\|^{q}_{q,V}}{V(x)}\leq\rho_{q,V}^{s}.$ (13)
Under this assumption the coefficients $\beta(s)$ and $\beta_{q}(s)$ can be bounded as follows,
[…]
thus the average-mixing time satisfies,
$t^{\sharp}_{\mathsf{mix}}(\xi)\leq\cfrac{\log\left\|\pi\right\|_{q,V}^{q}+\log\frac{1}{2\xi}}{\log\rho_{q,V}^{-1}},$
As a result, writing for simplicity, $\Pi_{r,V}\doteq\left\|\pi\right\|_{r,V}^{r}$ for $r\in(0,1)$, for any $q\in(0,(1-1/p)/2]$ we can bound the entropic term as follows,
$\begin{split}\mathcal{J}_{p}^{(s)}\leq\left(\rho_{q,V}^{s}\Pi_{q,V}+\Pi_{(1-1/p)/2,1}^{2}\right)^{2},\end{split}$
leading to the following theorem.
Theorem 4.1 (Under $V_{q}$-geometrical ergodicity).
Let $p\geq 1$. If there exists $q\in(0,(1-1/p)/2]$ such $P$ is $V_{q}$-geometrically ergodic—refer to (13)—for any $s\in\mathbb{N}$, it holds that
$\begin{split}\mathcal{J}_{p}^{(s)}\leq\left(\rho_{q,V}^{s}\Pi_{q,V}+\Pi_{(1-1/p)/2,1}^{2}\right)^{2}.\end{split}$
Observe from the above that, as soon as
[…]
we obtain a control of the entropic term in terms of a purely stationary quantity
$\begin{split}\mathcal{J}_{p}^{(s)}\leq 4\Pi_{(1-1/p)/2,1}^{4},\end{split}$
giving us insight into the exit time of the transient regime illustrated in Figure 3. Finally, a similar proof allows us to analyze the entropic term under the following condition, which is much weaker than $V$-geometrical ergodicity.
Theorem 4.2.
Suppose that there exist $\rho\in(0,1)$ and $V\colon\mathcal{X}\to[1,\infty)$ such that for any $s\in\mathbb{R}$,
$\sup_{x,x^{\prime}}\left|P^{s}(x,x^{\prime})-\pi(x^{\prime})\right|\frac{V(x^{\prime})}{V(x)}\leq\rho^{s}.$
Then for all $p\geq 1$ and for all $s\in\mathbb{N}$, it holds that
$\begin{split}\mathcal{J}_{p}^{(s)}\leq\left(\rho^{s(1-1/p)/2}\left\|\pi\odot V\right\|_{(1-1/p)/2}^{(1-1/p)/2}\left\|1/V\right\|_{(1-1/p)/2}^{(1-1/p)/2}+\Pi_{(1-1/p)/2,1}^{2}\right)^{2}.\end{split}$
@@
Obvious summability conditions on $\left\|1/V\right\|_{(1-1/p)/2}$ and $\left\|\pi\odot V\right\|_{(1-1/p)/2}$ are necessary for the bound in the above theorem to be non-vacuous. Additionally, under the condition of Theorem 4.2 a bound on the $\beta$-mixing coefficient can be immediately obtained as follows,
$\beta(s)\leq\frac{\rho^{s}}{2}\left\|\pi\odot V\right\|_{1}\left\|1/V\right\|_{1}.$